TREX: Logbook Entries

TREX: radiation Messages: 25 Entries..

Return to Logbook Contents Page
Entry Date Title Site Author #Graphics
249 Wed 26-Apr-2006Radiation Sensor Cleaningallagee
216 Thu 20-Apr-2006Rad.cleanedsouthmilitzer
212 Wed 19-Apr-2006Rad.cleanedcentralmilitzer
209 Tue 18-Apr-2006Rad.cleaned, Rnet leveledwestmilitzer
194 Sat 15-Apr-200610m Rad cleaned, photoscentralpoulos
177 Wed 12-Apr-2006'cleaned' 10m K&Zcentralmilitzer
166 Sun 09-Apr-2006Radiometers 'Cleaned'westmilitzer
165 Sun 09-Apr-2006Radiometers 'Cleaned'centralmilitzer
164 Sun 09-Apr-2006Radiometers 'Cleaned'southmilitzer
150 Tue 04-Apr-2006ZigBee now power on/off at Centralcentralsemmer
146 Tue 04-Apr-2006loggers at West and Southsouthsemmer
144 Mon 03-Apr-2006Intermittent logger datasouthhorst
135 Sun 02-Apr-2006Rad system ar west runningsemmer
131 Fri 31-Mar-2006logger at westwestsemmer
123 Wed 29-Mar-2006Logger #1 at centralcentralsemmer
121 Wed 29-Mar-2006Shutdown pyg sensor at central, logger ID 2centralsemmer
113 Mon 27-Mar-2006Battery changed at logger #1 at centralcentralsemmer
98 Thu 23-Mar-2006maintenance at west sitewesthorst
97 Thu 23-Mar-2006maintenance at central sitecentralhorst
96 Thu 23-Mar-2006Maintenance at south yesterdaysouthhorst
66 Fri 17-Mar-2006single pyg unitscentralsemmer
50 Tue 14-Mar-2006Tcase working at centralcentraloncley
47 Mon wiring errorcentraloncley
39 Sat 11-Mar-2006visitcentraloncley
36 Fri 10-Mar-2006Radiometer Power Issuescentralagee

249: radiation, Site all, Wed 26-Apr-2006 15:11:20 PDT, Radiation Sensor Cleaning
All of the radiation sensors were cleaned today.  South was cleaned at 12:25PST, West at 2:25PST, and Central at 2:43, 2:44, and 2:45 from south to north, respectively.  The net rad (1m) at central was particularly dirty to the point that it was difficult to see through the dome.  Unable to clean the 10m due to being at the site alone.  

216: radiation, Site south, Thu 20-Apr-2006 19:13:27 PDT, Rad.cleaned
20Apr06, 17:55

Cleaned all domes.  No real evidence of dirt on tissues or visually on domes.

212: radiation, Site central, Wed 19-Apr-2006 13:33:31 PDT, Rad.cleaned
19Apr06, 12:25-33

Cleaned all domes.  The Rnet and some of the long, short wave radiometers did show a bit of dirt on the tissue.  The Rnet had visible dust on its domes but not heavy.

209: radiation, Site west, Tue 18-Apr-2006 17:43:37 PDT, Rad.cleaned, Rnet leveled
18Apr06, 15:50-15:55

Cleaned all domes.  Absolutely no evidence of dirt.

Adjusted level on Rnet.
tilted it 'up' from a sag to the south that had the bubble outside
of the target but not pegged.
194: radiation, Site central, Sat 15-Apr-2006 12:44:09 PDT, 10m Rad cleaned, photos
15Apr06, 11:00PDT
Cleaned 10m Rad.  No visible evidence of dust.
A full set of 360, etc photos were taken from the 10m perspective.

177: radiation, Site central, Wed 12-Apr-2006 11:33:03 PDT, 'cleaned' 10m K&Z
12Apr06, ~10:15
attempted to clean the radiometer with methanol using Steve Semmer's sponge mop.
The wind was substantial so I only got the downward looking dome swiped.  There was no evidence of any dirt either by visual inspection or by looking at the 'swabber.'  Consequently I did not maneauver around the logger, etc. to swipe the upper looking dome.   The suspicion is there is just too much wind for this very fine dust to stick, and not enough moisture.   We have not seen any rain.
166: radiation, Site west, Sun 09-Apr-2006 17:45:35 PDT, Radiometers 'Cleaned'
16:30PDT 4/9/06 Cleaned radiometer domes.
These were all in good shape and did not appear to have any 'dirt' on them.

165: radiation, Site central, Sun 09-Apr-2006 14:33:50 PDT, Radiometers 'Cleaned'
11:30PDT 4/9/06 Cleaned radiometer domes.
These were all in good shape and did not appear to have any 'dirt' on them.
Note: Did not clean 10m rad

164: radiation, Site south, Sun 09-Apr-2006 12:42:28 PDT, Radiometers 'Cleaned'
10:00PDT 4/9/06 Cleaned radiometer domes.
These were all in good shape and did not appear to have any 'dirt' on them.
Note: slightly adjusted Rnet level
150: radiation, Site central, Tue 04-Apr-2006 16:09:33 PDT, ZigBee now power on/off at Central
The two single pygreometer logger boxes, #1 and #2, now have their radios
being turned on then off by the logger. They are setup to be on for 1 second
and off for 4 seconds. New batteries were insatlled in each unit. The
code change can be found on the Panasonic under the PC208W directory.
The file name is TREXzig.

146: radiation, Site south, Tue 04-Apr-2006 08:51:55 PDT, loggers at West and South
I changed the transmit power of the ZigBee units at west and south. They are now down to 10 dBm verses 18dBm. 
144: radiation, Site south, Mon 03-Apr-2006 15:50:26 PDT, Intermittent logger data
There is once again intermittent logger data showing up in the netcdf files and
Splus plots.  However, I have scanned an entire raw data file for a time when 
this occurs and see not missing data.
135: radiation, Site , Sun 02-Apr-2006 08:00:17 PDT, Rad system ar west running
Tom notice that the 2.4 Ghz radios did not solve the missing radiation data.
So after getting the hot film going, we took the extension serial cables for
the hot film sonics to hardwaire in the logger at west. This took place around
16:30 yesterday.
131: radiation, Site west, Fri 31-Mar-2006 12:26:46 PST, logger at west
The Campbell data loggers at south and west have been losing data intermittently
which Steve suspects may be due to 915 MHz interference.
He removed the logger at west around 11:45 PST to replace the MaxStream radio 
with a Zigbee radio.

   Went to re-install logger at west at about 16:00. For some unknown reason the cr10 lost its program. Will fix tomorrow.

Update 4/1/06:  Steve went to reprogram logger and after switching cables to do this, it started on its own, so apparently did not lose its program as thought.

123: radiation, Site central, Wed 29-Mar-2006 16:53:13 PST, Logger #1 at central
The logger # 1 at central is back up (~16:15). The code changes to turn on
and off the ZigBee did not work. The transceiver comes on but the receive
light does not. The logger was put back into tis old mode.

121: radiation, Site central, Wed 29-Mar-2006 10:18:24 PST, Shutdown pyg sensor at central, logger ID 2
I have removed logger ID 2 from central to make some modes to the software
so that the radio will go to sleep between data transmissions. This should
help the power issue at the two single pygreometer sites, The logger was
removed around 9:15 am local,

113: radiation, Site central, Mon 27-Mar-2006 11:42:15 PST, Battery changed at logger #1 at central
The battery was changed at radiaiotn logger #1 at central.
This took place around 10:45 this morning.

98: radiation, Site west, Thu 23-Mar-2006 15:43:15 PST, maintenance at west site
March 23: Tom at west site for maintenance; arrived ~ 1415 pst

Guywire tensions (lbs)
1	450	360	375
2	300	290	310
3	385	300	280
4	340	295	290

Cleaned radiometers with methanol 14:38
Replace fuse on serial channel 14 to restart aerosol spectrometer.
Alone, so did not clean KH2O
97: radiation, Site central, Thu 23-Mar-2006 13:47:19 PST, maintenance at central site
March 23: Steve and Tom
Service visit to central site
Guywire tensions:
1	380	450	420
2	325	310	350
3	310	370	300
4	320	300	305

Steve cleaned KH2O and LI-7500 with methanol
Steve on 30 m tower 11:35 - 12:05
Steve of 5 m tower  12:05 - 12:10

Then cleaned radiometers with methanol spray, wiped off radiometers
with kleenex.

96: radiation, Site south, Thu 23-Mar-2006 10:38:56 PST, Maintenance at south yesterday
Steve and Tom visited south tower Mar 22 before noon for routine maintenance.
Measured guywire tensions, all above 300 lbs:

1	400	370	410
2	360	325	320
3	400	320	340
4	340	330	315

Steve cleaned radiometer domes and KH2O windows with methanol.
He wiped off the net radiometer domes.
All radiometers appeared fairly clean prior to cleaning.

We crashed NIDS by running rserial locally.  This is a known
bug.  Should have run through ASTER (ssh aster@server)

66: radiation, Site central, Fri 17-Mar-2006 10:55:58 PST, single pyg units
Spent some time out at central changing the batteries and re-orienting the solar panels at the two single pyg sites. This took place between 9:00 and 10:30 am.

50: radiation, Site central, Tue 14-Mar-2006 11:57:26 PST, Tcase working at central
After several conversations, Steve S/John/I found a programming error in the
central 4-comp logger.  The P7 instruction needed single-ended channel 3, not
2.  Made the change manually (*7h*1...*0) and cycled power to check that it
was saved.  Now Tcase ratios are 1.35, 1.34, and 1.34 for a1, a2, and 4-comp
"out"s, respectively!  The 10m values are different, but perhaps this is 
real and/or a calibration issue.

47: radiation, Site central, Mon 13-Mar-2006 12:21:25 PST, wiring error
Found that the wire logger L2 to Amp pin 6 had been inserted in pin 8 by 
mistake.  Fixed this about 11:30.  Didn't seem to help the Tcase data :(

39: radiation, Site central, Sat 11-Mar-2006 11:12:54 PST, visit
Was at central ~10:15 to:
- check fuel cell: 13.45V/1260psi
- clean all rads (including 10m for the first time)
- TRIME: 38%/11%/28% (from west to east) -- basically unusable.
- check logger cabling.  Only difference between a2 and 4-comp on K&Z is white
wire goes to "AG" on a2 rather than "G" on 4-comp.  TP01 wiring looks the
same as I remember at west.

36: radiation, Site central, Fri 10-Mar-2006 11:01:39 PST, Radiometer Power Issues
Steve and I spent this morning trying to figure out why two of the radiation stands at central and the rad stand at west were not working.  A simple fix to get them running while we troubleshoot the problem was to switch out the batteries with ones we had in the trailer.  There was a decent amount of dust on the solar panels at central, but we are unsure how much of a factor it is.